A bottle maker believes that 12% of his bottles are defective.

If the bottle maker is accurate, what is the probability that the proportion of defective bottles in a sample of 412 bottles would be greater than 11%? Round your answer to four decimal places.
1
Expert's answer
2020-10-15T17:45:55-0400

solution


Population proportion, "p=0.12"

Sample proportion, "\hat p=0.11"



"Z = \frac{\hat p -p}{\sqrt{\frac{p(1-p)}{n}}}"

"\frac{0.11 -0.12}{\sqrt{\frac{0.12(1-0.12)}{412}}}=-0.6246"

When the Z score is -0.6246, the probability that "Z>-0.6246" is "0.73388"


answer:

The probability that the proportion is greater than 0.11 is 0.7339